i need to find how how to do factoring
5t2 - 10t = 0

Found 2 solutions by checkley77, smalltimegenius:
Answer by checkley77(12844)

5t^2 - 10t = 0
5t(t-2)=0
t-2=0
t=2 answer.

Answer by smalltimegenius(2)